OpenStudy (anonymous):

determine alegebraically whether the function is even, odd or neither f(x) = x/(x^2+5) I orginally assumed it was even because of x^2

hartnn (hartnn):

Find f(-x) by replacing x in f(x) by -x. If f(-x) = f(x), then the function is even If f(-x) = -f(x), then the function is odd else neither.

OpenStudy (anonymous):

how would you write/solve for -f(x)?

hartnn (hartnn):

-f(x) is just -x/(x^2+5)

OpenStudy (anonymous):

so... would it be\[f(x) = - (x \div x^{2} + 5)\] ??

hartnn (hartnn):

to get f(-x) you just replace x by -x in the f(x) equation what u get for f(-x) then ?

OpenStudy (anonymous):

He is saying solve f(-x) so it would be \[f(-x)=\frac{ (-x) }{ (-x)^2+5 }\]If f(-x) = f(x), then the function is even If f(-x) = -f(x), then the function is odd else neither.

OpenStudy (anonymous):

\[f(-x) = -x \div x^{2}+5\] because x^2 cancels out the negative but what I dont understand is why when it's -f(x), shouldn't the 5 be negative too? therefore it would be "neither"?

OpenStudy (anonymous):

because then it would be\[f(-x) = -x \div x^{2} + 5 and -f(x) = -x \div x^{2}-5\]

hartnn (hartnn):

no, here u get f(-x) exactly equal to -f(x) so its odd and you only replace x by -x so 65 will remain +5

hartnn (hartnn):

sorrry, +5 will remain +5

OpenStudy (anonymous):

I thought -f(x) made the whole equationo negative

hartnn (hartnn):

yes, whole means only numerator... so x in numerator became -x

OpenStudy (anonymous):

oooo, I had no idea.. I hate overthinking these things -.-

hartnn (hartnn):

so now its clear, how its odd ?

OpenStudy (anonymous):

yep!